OpenStudy (anonymous):

Simplify f(x) = 2x^2+3x+1/2x+1 Using complete sentences, explain why f(1) = 2, f(0) = 1, and f(-1) = 0, yet f(-1/2) is undefined. Make sure to show your work

OpenStudy (anonymous):

lol i always laugh when i read a question that asks to write in sentences for math xD

OpenStudy (anonymous):

u there? @forever.grace

OpenStudy (anonymous):

why do you laugh?

OpenStudy (anonymous):

cause math is about calculating and not right sentences lol

OpenStudy (anonymous):

oh very true..

OpenStudy (anonymous):

well to ur question :)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

lol u wanted help with the post question right? isnt that y u post it ;)

OpenStudy (anonymous):

oh yea lol.

OpenStudy (anonymous):

haha it's ok im fun to talk to :P

OpenStudy (anonymous):

f(x) = 2x^2+3x+1/2x+1

OpenStudy (anonymous):

lol okay?

OpenStudy (anonymous):

f(x) = 2x^2+3x+1/2x+1 f(x) = (2x + 1)(x +1)/2x+1 f(x) = x +1 now isnt that easier to look at :P

OpenStudy (anonymous):

understand wat i did? :)

OpenStudy (anonymous):

yea i think so.

OpenStudy (anonymous):

2x^2+3x+1 = ( 2x + 1 ) ( x + 1 ) and since \[\frac{ ( 2x + 1 ) ( x + 1 )}{ 2x+1 }\]

OpenStudy (anonymous):

the 2x + 1 cancel out and ur left with x+1, is that better?

OpenStudy (anonymous):

much better.

OpenStudy (anonymous):

cool :)

OpenStudy (anonymous):

therefore f(x) = x + 1

OpenStudy (anonymous):

f(1) = (1) + 1 f(1) = 2

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

well we havnt finished the question yet so dont thank me yet :P

OpenStudy (anonymous):

ohh lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

also u need to put everything i just did in words :P

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now u show that f(0) =1

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

f(-1/2) is undefined because you can't divide by zero and 2x+1 is zero for x=-1/2.